OpenStudy (anonymous):

find the product of -4(v+1)

OpenStudy (anonymous):

-4v-4

OpenStudy (anonymous):

multiply -4 times V then multiply it by 1. which is -4V + 1

OpenStudy (anonymous):

then how do you get the other -4? that NK309

OpenStudy (anonymous):

got

OpenStudy (anonymous):

no its the distributive property. a (b+c) =ab + ac

OpenStudy (anonymous):

mmm, what? lol

OpenStudy (anonymous):

if you ever have a problem like that, # (#+#), use the distributive prop

OpenStudy (anonymous):

i dont know how. :o

OpenStudy (anonymous):

to 'use the distributive property' is the same as 'multiply it out, or expand the brackets' so #1(#2 + #3) = (#1 times #2) plus (#1 times #3)

OpenStudy (anonymous):

so like.. v+1=1v then x-4? or am i totally wrong?

OpenStudy (anonymous):

wrong, but you seem to be getting closer :) v+1 is different to 1v because 1v is one multiplied by v just like v+4 is different to 4v because instead of having 4 + v (or 4 + 1v) its 4 multiplied by v does that make sense?

OpenStudy (anonymous):

kindaaa. but i dont get how they get -4x-4

OpenStudy (anonymous):

i got -4x+-8

OpenStudy (anonymous):

its because of the minus sign out the left side of the 4 -4 (v+1) ^ the one above this: ^ that minus sign makes it negative 4 multiplied by each thing in the bracket, so you have to change the sign every time

OpenStudy (anonymous):

you could get -4x+-4 but if you got -4x+-8 then you would have had -4(x+2)

OpenStudy (anonymous):

lol.. what the hell am i doing. :/

OpenStudy (anonymous):

you were so close that time.... i reckon you'll have it ont he next problem

OpenStudy (anonymous):

maybe, lol.

OpenStudy (anonymous):

you can do it!